Tonight (Oct 31), certain NJIT websites will be unavailable due to critical security upgrades between 10pm and 12am. The list of websites include: Information Services and Technology Center for Wireless Communications and Signal Processing Enterprise Development Center Office of Strategic Communications Admin Council Staff Council Graduate Studies Institutional Research and Planning Shared Governance These websites andContinue reading “Website and server maintenance – Oct 31 10pm-12am”
The primary NJIT.edu webserver will be offline for approximately 5-10 minutes while services are being upgraded at 12:00am, Saturday, May 3. During this time, the majority of all NJIT websites may be unresponsive until the upgrade is complete. For additional information or questions, please contact firstname.lastname@example.org.
An AFS fileserver will be unavailable from about 7:30 am to 8:00 am today for an emergency hardware repair. The login directories of some AFS users, as well as some Web pages and research directories, will be unavailable during this time. Affected users logging in to an AFS client will be put in the directoryContinue reading “Unavailability of Some AFS Login & Other Directories”
Campus Pipeline service has been restored as of 8:45 am. We will continue to monitor it.
Campus Pipeline is currently unavailable. We are researching the problem. We will update you later when we have more info.
As of 5:00 am this morning, Oracle engineers have made the disk storage array available for NJIT applications and services. At this time, ADM email and related services are available. (ADM email (aka Outlook email) is typically used by NJIT faculty and staff.) Other services will be “re-started” and made available to the campus community. Continue reading “Status update at 8:00 am”
As of 6:00 pm this evening, Oracle engineers have successfully reconstructed all of the failed components of the disk storage array that has been out of service since early Tuesday morning. They are now in the process of reconnecting these components to make all of them available to the storage array’s controllers. This process willContinue reading “Status update at 8:00 pm”
As of noon today, the process to export 23 reconstructed volume groups was completed, a failed disk controller was replaced and now dual controllers in the storage array are functioning normally. Five of 18 volume groups needing to be reconstructed have been successfully reconstructed. Recovery procedures are run with due caution and continued success byContinue reading “Major service interruption update at 12:30 pm”
At midnight this morning, all automated data reconstruction was completed successfully on 23 of 41 volume groups comprising the disk array. The remaining 18 of 41 volume groups will be reconstructed in a less automated fashion, as was expected. A process to export the 23 reconstructed volume groups is nearly completed. This process serves toContinue reading “Major service interruption update 7:15 am”
Normal Moodle services are now available. Disk array data recovery is proceeding successfully, but very slowly. This recovery process is being closely monitored. As soon as possible after it is completed, other services will be restored to their normal working conditions. Your continued understanding is appreciated.